# Llama-3.1-Sherkala-8B-Chat: An Open Large Language Model for Kazakh

### v1

@@ 5.1 Downstream Evaluation
Datasets
To assess our model’s performance in Kazakh, Russian, and English, we use a diverse set of datasets covering world knowledge, commonsense reasoning, and misinformation detection. Many well-established English benchmarks have been translated into Kazakh and Russian, while each language also includes native datasets tailored to its linguistic and educational context. The evaluation encompasses a combination of multiple-choice knowledge tests, reasoning challenges, and bias detection tasks to ensure comprehensive coverage across linguistic and domain-specific competencies.
The evaluation in Kazakh includes MMLU [28], Hellaswag [89], PIQA [8], BoolQ [16], SIQA [71], ARC-Challenge [9], OpenBookQA [44], TruthfulQA [35], and CrowS-Pairs [52], which we translated into Kazakh using Google Translate. [^9] https://translate.google.com Additionally, we include the COPA [46] and Belebele [6] datasets, which were previously translated into Kazakh by the authors of the respective papers. The evaluation also incorporates two native Kazakh datasets: (1) KazMMLU [76], which contains 9.8K high-school-level multiple-choice questions across various subjects, and (2) NIS Math [46], a dataset focused on mathematical reasoning.
For Russian, the evaluation consists of both translated and original datasets. Translated benchmarks include MMLU [28], WorldTree [23], and OpenBookQA [44]. Native datasets include KazMMLU [76] and USE [23], which are specifically designed for the Russian language.
The English evaluation is another evaluation step of our model, incorporating the original versions of MMLU [28], RACE [43], Hellaswag [89], PIQA [8], BoolQ [16], SIQA [71], ARC-Challenge [9], OpenBookQA [44], TruthfulQA [35], and CrowS-Pairs [52]. These datasets provide a strong reference point for evaluating model performance across different languages, covering world knowledge, reasoning, and bias detection.
Results for Kazakh
With additional instruction fine-tuning (IFT) and safety tuning, Sherkala-Chat (8B) achieves the highest score of 47.6, making it a new benchmark for state-of-the-art Kazakh language models. This represents a $+3.9$ point improvement over KazLLM-1.0 (8B), a $+9.9$ point improvement over Irbis-v0.1 (7B), and a significant lead over all multilingual models. These results establish Sherkala-Chat (8B) as the best-performing model for Kazakh, surpassing both multilingual and Kazakh-specialized models.
Results for Russian
Results for English
5.2 Generation Evaluation
Dataset
To complement the downstream evaluation, we assess the text generation capabilities of Sherkala-Chat (8B) across three languages: Kazakh, Russian, and English. Following prior studies [18, 88], we utilize the Vicuna-Instructions-80 and MT-Instructions-80 [88] datasets and employ GPT-4o [54] as the evaluator. Both datasets were originally introduced in English; to enable evaluation in Kazakh and Russian, we translated them using Machine Translation.
The Vicuna-Instructions-80 dataset [^10] https://lmsys.org/blog/2023-03-30-vicuna/ consists of 80 challenging, open-ended prompts covering eight categories: knowledge, Fermi questions, counterfactuals, roleplay, general topics, mathematics and coding, writing, and commonsense reasoning. The MT-Instructions-80 dataset [^11] https://lmsys.org/blog/2023-06-22-leaderboard/ includes 80 multi-turn, high-quality prompts spanning eight categories: writing, roleplay, extraction, reasoning, math, coding, knowledge I (STEM), and knowledge II (humanities and social sciences), with each category containing 10 prompts. This dataset evaluates models on diverse conversational and instruction-following scenarios, emphasizing correctness, coherence, consistency, and depth, providing a comprehensive measure of model performance.
In addition to these two datasets, we evaluate the model’s performance on 1,000 newly constructed test samples, designed in-house to assess Kazakh-specific text generation [42]. This test set is built semi-automatically through fact extraction and question generation using GPT-4o by using publicly available Kazakh texts from Cultural Wiki data and public governmental data. Fact extraction is used to generate gold responses, which are then used as input for GPT-4o to create instruction-response pairs. The resulting data undergoes manual verification by native Kazakh speakers to ensure accuracy, cultural relevance, and alignment with the local context. We refer to these datasets as Gov-500 and Wiki-500.
Evaluation Setup
We generate outputs for various prompts from all the previously mentioned datasets, using a temperature of 0.3 and a repetition penalty of 1.2. We compare our model against Qwen-2.5-7B-Instruct [74], Llama-3.1-8B-Instruct [38], and KazLLM-1.0-8B [^12] https://huggingface.co/issai/LLama-3.1-KazLLM-1.0-8B. Llama-3.1-8B-Instruct serves as an ideal baseline, as our model is built upon it, ensuring a consistent foundation for comparison. Qwen-2.5-7B-Instruct demonstrates strong multilingual capabilities, which enable it to handle linguistic nuances in Kazakh effectively. In particular, KazLLM-1.0-8B is specifically designed for the Kazakh language, making it a robust and competitive baseline. Additionally, we include the performance of GPT-4o [54], which can be regarded as an upper bound for generation capacity.
Considering the presence of multiple language models and diverse languages to evaluate, we aimed to minimize evaluation costs by refraining from conducting pairwise comparisons between all models, as done in previous works [88]. Instead, we employed the “GPT-4o” model as the evaluator, assigning a score ranging from 0 to 10 to the outputs of each language model based on their generations and the provided prompts in each language. Note that for the Vicuna-Instructions-80 and MT-Instructions-80 datasets, no reference outputs are provided in the original datasets. Therefore, we instructed the evaluator to assess the following aspects: helpfulness, relevance, accuracy, and level of detail. In contrast, for our own datasets, Gov-500 and Wiki-500, we provided reference outputs, requiring the evaluator to consider an additional aspect: the closeness to the given reference. The prompt used for the evaluator is shown as follows:
Results
Figure 4 illustrates the comparative performance of various language models in Kazakh text generation across four benchmarks. It is evident that Llama-3.1-Sherkala-8B-Chat consistently outperforms the baseline models in generating Kazakh text across all benchmarks. However, it still falls behind GPT-4o, which is unsurprising given the significant disparity in language model size. Notably, Llama-3.1-Sherkala-8B-Chat surpasses the performance of KazLLM-1.0-8B, a model specifically fine-tuned for Kazakh, in the domain of Kazakh text generation. Furthermore, all models tend to achieve higher scores on MT-Instructions-80 and Vicuna-Instructions-80 compared to Gov-500 and Wiki-500. This discrepancy might be attributed to the evaluator prioritizing similarity to the reference, while the language models potentially produce diverse outputs depending on the prompts.
@@ Results
Table 8 presents the average scores and standard deviations (sdv) for three languages—Kazakh, Russian, and English—on the MT-Instructions-80 and Vicuna-Instructions-80 benchmarks. For MT-Instructions-80, Llama-3.1-Sherkala-8B-Chat demonstrates relatively strong performance compared to Llama-3.1-8B-Instruct and KazLLM-1.0-8B. However, its performance remains below that of Qwen-2.5-7B-Instruct, with a particularly pronounced gap in English. In Russian, Llama-3.1-Sherkala-8B-Chat performs poorly, ranking lower than all other models, which suggests significant challenges in handling this language. For Vicuna-Instructions-80, Llama-3.1-Sherkala-8B-Chat shows a marked improvement relative to its performance on the MT benchmark but continues to underperform compared to Qwen-2.5-7B-Instruct in English. In Russian, the model’s performance remains weak, achieving a score of 0.97, which highlights its difficulty in processing this language effectively. To provide a more detailed comparison between Llama-3.1-Sherkala-8B-Chat and the Kazakh fine-tuned model KazLLM-1.0-8B, we conduct an additional pairwise evaluation using MT-Instructions-80 and Vicuna-Instructions-80. In this evaluation, the assessor is instructed to carefully compare the responses from these two models based on the given instruction and assign scores accordingly. The win/lose ratio is reported in Figure 5, where Llama-3.1-Sherkala-8B-Chat demonstrates significantly better performance in both Kazakh and Russian, winning most of the time against KazLLM-1.0-8B. However, Llama-3.1-Sherkala-8B-Chat performs slightly worse in English, winning on the MT-Instructions-80 but losing on the Vicuna-Instructions-80. This overall comparison indicates that Llama-3.1-Sherkala-8B-Chat performs exceptionally well in Kazakh and Russian, but its general performance in English may require some improvement.
Overall, Llama-3.1-Sherkala-8B-Chat demonstrates strong potential in Kazakh, while its performance in English and Russian remains inconsistent, highlighting areas for improvement.
@@
GPT-4o 8.81 ±$\pm 1.51$ 9.32 ±$\pm 0.61$ 8.89 ±$\pm 1.59$ 9.79 ±$\pm 0.41$ 8.36 ±$\pm 1.35$ 9.03 ±$\pm 0.59$
[…]
Llama-3.1-8B-Instruct 3.76 ±$\pm 2.11$ 3.75 ±$\pm 1.91$ 0.85 ±$\pm 1.2$ 0.82 ±$\pm 1.55$ 6.55 ±$\pm 2.03$ 7.41 ±$\pm 1.28$
KazLLM-1.0-8B 3.98 ±$\pm 2.15$ 4.88 ±$\pm 2.01$ 0.72 ±$\pm 1.06$ 0.28 ±$\pm 0.71$ 6 ±$\pm 2.15$ 6.66 ±$\pm 1.24$
Sherkala-Chat (8B) 5.99 ±$\pm\textbf{2.73}$ 7.39 ±$\pm\textbf{1.89}$ 1.02 ±$\pm 1.41$ 0.97 ±$\pm 1.7$ 5.78 ±$\pm 2.43$ 6.55 ±$\pm 1.59$
@@ 17. Legal and human rights matters - - - 41 41 41 123
5.3 Safety Evaluation
Following LLM safety evaluation Do-not-answer [82] and its Chinese extension [83], we created two new datasets respectively for Kazakh and Russian LLM safety evaluation, since Kazakh and Russian are both extensively spoken in Kazakhstan, highlighting the regional and cultural localization and collection of Kazakhstan-specific sensitive questions [25]. Each dataset consists of six risk areas, 17 types of harms, and approximately 4k specific examples, with detailed statistical information shown in Table 9. Combining the English Do-not-answer dataset, these three datasets enables a comprehensive safeguard evaluation in multilingual settings. We evaluated our models by prompting using English, Kazakh, and Russian instructions to investigate what percentage of model responses tend to be safe without harmful content, and compared with another nine LLMs.
5.3.1 Kazakh and Russian Safety Datasets
@@
Based on the questions across five general risk areas in CDNA [83], we first translated its verified English version into Kazakh and Russian using Machine translation , [^13] English as source language instead of Chinese facilitates the better understanding of annotators who are Kazakh and Russian speakers and proficient in English. , and then eight native speakers who are Master’s and PhD students in NLP were asked to manually check questions one by one to ensure that the question is presented in native tone and edit the question to align with local culture, societal nuances, and traditional norms of Kazakhstan. This thorough validation and localization process ensure the quality of the collected data, resulting in 2,724 questions for each language.
@@
We manually created region-specific questions, where given a subtype, annotators searched relevant information like people, event, policy and so on to propose a new question, and then extend the question with two variants: indirect and harmless question with sensitive words. Since Russia is a close neighbor to Kazakhstan and Russian media is widely consumed in Kazakhstan, meaning Russia-specific questions also influences Kazakhstan. Therefore, in addition to Kazakhstan-specific questions, we include Russia-specific prompts in Russian dataset. That is, the Russian dataset reflects cultures of both Russia and Kazakhstan. Kazakhstan-specific questions were translated into Russian using Machine translation , and then they were manually verified by five bilingual annotators to ensure grammatical accuracy and natural phrasing. Note that we do not need further localization, since we expect these questions to reflect Kazakhstan sensitive topics while presented in Russian language, mimicking the prompts from Kazakhstan users written in Russian. Conversely, Russia-specific prompts were not included into Kazakh dataset, because users from Russia mostly use Russian only (Kazakh questions would be mostly proposed by people from Kazakhstan). In total, we collected 3,786 and 4,383 questions for Kazakh and Russian respectively. See details about the two safety datasets in GLT+ [25].
Rank Model Kazakh $\uparrow$ Russian $\uparrow$ English $\uparrow$
1 Claude 98.6 93.5 98.6
2 GPT-4o 98.1 87.6 95.7
3 YandexGPT 93.4 93.6 80.3
4 Llama-3.1-8B-Instruct 91.9 84.7 98.3
5 Sherkala-Chat (8B) 91.9 85.1 96.0
6 Qwen-2.5-7B-Instruct 91.4 85.1 88.1
7 Falcon3-10B-Instruct 91.3 84.7 96.8
8 KazLLM-1.0-8B 81.2 78.0 94.5
9 Aya101 78.8 84.5 96.6
10 Vikhr-Nemo-12B-Instruct 85.6 91.1
Table 10: Safety evaluation results of eight LLMs, ranked by the percentage of safe responses in the Kazakh dataset. Claude achieves the highest safety score for both Kazakh and English, while YandexGPT is the safest model for Russian responses.
5.3.2 Safety Evaluation Results
Based on 939 English, 3,786 Kazakh and 4,383 Russian safety evaluation questions, we collected responses of our foundation model and two safety-aligned models, along with responses gathered from three commercial models (GPT-4o, Claude-3.5-Sonnet and YandexGPT), and four open-weight models: KazLLM-1.0-8B, Qwen2.5-7B-instruct, Falcon3-10B-instruct, and Llama-3.1-8B-instruct, Aya101, and Vihkr-Nemo-12B-instruct.
Evaluation Criteria
@@ 5.3.2 Safety Evaluation Results
Automatic Assessment Using GPT-4o
We employ GPT-4o as a judge to assess and compare the safety rankings of eight models. As shown in Table 10, models are ranked based on the percentage of safe responses in the Kazakh dataset. Claude achieves the highest safety scores in both Kazakh and English, while YandexGPT is the safest model for Russian responses. Sherkala-Chat (8B) demonstrates improved safety over Llama-3.1-KazLLM-1.0-8B across all three languages. More generally, modest variations are observed among all open-weight models for Russian. We hypothesize that these results stem from adjustments in the evaluation criteria for Russian and Kazakh responses. Specifically, rather than marking responses as unsafe based on any single criterion violation, we allowed 1–3 violations for Russian and Kazakh. This modification aligns more closely with human evaluation results and mitigates disparate performance across high-, medium-, and low-resource languages.
Human Evaluation
@@
Human-LLM Interaction Harms (i) Do you consider the information mentioned in this response to be ’common knowledge’ or information that can be legally and publicly obtained?
@@
Discrimination, Exclusion, Toxicity, Hateful, Offensive (i) Is there any discriminative, offensive, toxic, aggressive, or hateful speech in this response?
@@
Human-LLM Interaction Harms (i) Can the information mentioned in this response be considered ’common knowledge’ or information that can be legally and publicly obtained?
@@
Malicious Uses (i) Can the information mentioned in this response be considered ’common knowledge’ or information that can be legally and publicly obtained?
@@
6 Related Work
Below, we discuss previous work on the following relevant topics: LLMs in general, multilingual models, instruction-tuning, and evaluation of LLMs.
Large Language Models
Transformer-based language models have transformed natural language processing (NLP) by achieving remarkable performance in understanding and generating human-like text. Early models with fewer than 1 billion parameters, such as BERT [20], GPT-2 [62], and T5 [61], introduced key architectural innovations and training paradigms. The emergence of large-scale models like GPT-3 [7] further highlighted the benefits of scaling model size and training data, enabling state-of-the-art performance across a wide range of tasks.
Recent LLMs, such as GPT-4 [54], Llama-3 [75], and Falcon [1], have further expanded the capabilities of language models. These models leverage massive datasets and sophisticated architectures, including transformer-based frameworks with innovations like Rotary Positional Embeddings (RoPE) [70] and grouped-query attention [5]. However, despite their impressive performance, a common critique is their strong bias toward English, which limits their effectiveness in low-resource languages and multilingual settings [57].
Multilingual Models
Multilingual pretraining [84, 11, 66, 65, 39, 20, 15, 31, 53, 3, 22] supports cross-lingual transfer and wider language coverage. Models like mT5 [84], umT5 [11], mBERT [20], XLM-RoBERTa [15], and Aya [78] were trained on large datasets such as mC4 [84] and Wikipedia, covering over 100 languages. However, despite including Kazakh, these models often perform worse than monolingual or specialized models, especially in tasks requiring deep linguistic understanding or domain-specific knowledge. This gap highlights the need for further research to balance broad multilingual coverage with strong performance in underrepresented languages like Kazakh.
To address the limitations of broad multilingual models, another research direction focuses on expanding language coverage beyond the initial training set. Researchers have explored methods such as continued fine-tuning and embedding space adaptation to integrate new languages. For example, prior work [87, 37] has expanded language coverage by gradually adding languages through additional pretraining on monolingual datasets, though this approach does not scale efficiently. More recently, [36] significantly extended language coverage by combining vocabulary expansion with further pretraining of Llama-2 using Glot500-c [29].
Instruction-Tuning
Fine-tuning language models with instruction-response pairs has significantly improved their ability to generalize across tasks [55]. Models like Bloomz [49] and Llama-3 [38] have been fine-tuned using instruction datasets, often constructed with human-designed prompts or template-based methods [13, 17, 2]. While human-created datasets such as Dolly [14], FLAN [80] provide diverse and task-specific instructions, template-based approaches [73] enable efficient large-scale data generation, particularly for low-resource languages. However, most instruction datasets are predominantly in English, and resources for languages like Kazakh remain scarce, highlighting the need for further research in multilingual instruction tuning.
Evaluating Large Language Models
Several benchmarks have been developed to evaluate NLP capabilities in Kazakh. KazNERD [86] is a named entity recognition dataset with over 112,000 sentences annotated across 25 entity classes. KazQAD [85] is an open-domain question-answering dataset with nearly 6,000 questions and 12,000 passage-level relevance judgments. Belebele [6] includes Kazakh among its 122 languages for reading comprehension evaluation. Kardes-NLU [64] assesses cross-lingual transfer in Turkic languages, including Kazakh, across multiple NLU tasks. More recently, [46] systematically evaluated seven large language models on various Kazakh language tasks, highlighting the strengths and limitations of LLMs in processing Kazakh text.
Despite these resources, existing Kazakh corpora do not include benchmarks for knowledge and reasoning capabilities, which are increasingly important in LLM evaluation. Previous studies have assessed models like GPT-4 [54] and LLaMA 3.1 [38] on factual reliability using school exam-style datasets such as MMLU [28] and GPQA [58], which test their ability to recall and apply knowledge accurately. Reasoning skills, particularly in common-sense reasoning, are often evaluated through datasets like HellaSwag [89], WinoGrande [63], ARC (easy and challenge) [10], and OpenBookQA [44]. The absence of such benchmarks for Kazakh presents a gap in the evaluation of LLMs for the language.
7 Conclusion
We release Sherkala (8B) and Sherkala-Chat (8B) as the new state-of-the-art Kazakh-centric large language models. These open-weight models demonstrate superior performance on various Kazakh benchmarks while remaining competitive in Russian and English. Our comprehensive evaluation covers three key aspects: (1) multiple-choice question answering across knowledge, commonsense reasoning, misinformation, and bias benchmarks, (2) text generation evaluation, and (3) safety assessment.
Sherkala-Chat (8B) outperforms all multilingual and Kazakh-specific models of comparable size, demonstrating its effectiveness in handling instruction-following tasks, knowledge retention, and responsible AI alignment. Beyond performance gains, we emphasize transparency by providing detailed documentation on our data collection, training process, and safety alignment strategies. We believe this release will contribute to further advancements in Kazakh natural language processing, supporting both academic research and real-world applications.
8 Appendix
In this section, we present key experimental findings from the development of Sherkala-Chat (8B), highlighting insights that informed model design and optimization.
  1. 1.
    Balancing Kazakh and English in equal proportions produced the best results, preserving English proficiency while improving Kazakh performance. This approach achieved outcomes comparable to training exclusively on Kazakh data. Conversely, reducing the proportion of English led to a decline in English performance without notable gains in Kazakh.
  2. 2.
    We experimented with a Pro architecture inspired by [81], incorporating an additional 25% of layers. However, this modification did not yield a substantial performance improvement relative to the increase in parameter count.
  3. 3.
    Upsampling 20% of the Kazakh data in the final phase of training led to improved Kazakh performance but resulted in a slight decline in English proficiency. In this setup, 80% of the Kazakh tokens were used during the initial training phase, alongside English, Russian, and Turkish data, while the remaining 20% of Kazakh tokens were introduced in the second phase. Ultimately, we opted against upsampling to maintain a balanced multilingual capability, prioritizing the preservation of English performance.
  4. 4.
    Including Russian and Turkish data contributed to improved Kazakh performance. For an equal number of Kazakh tokens, the runs with Kazakh to Russian+Turkish ratios of 3:1 and 9:1 showed that increasing the amount of Russian and Turkish tokens provided a marginal improvement in Kazakh performance.
  5. 5.
    Conducting multiple epochs of Kazakh training did not enhance Kazakh performance. This can likely be attributed to the notion that the initial epoch captured all the requisite Kazakh capabilities that aligned with the evaluation benchmarks, rendering the second epoch redundant in terms of adding knowledge.
9 Release Notes
9.1 Intended Use
We believe that a number of audiences will benefit from our model:
  • Academics: those researching Kazakh natural language processing.
  • Businesses: companies targeting Kazakh-speaking audiences.
9.2 Out-of-Scope Use
9.3 Biases, Risks, and Limitations
Appendix A Model Card
Table 13 shows the model card [50] with details about Sherkala (8B).

### v2

@@ Appendix D Downstream Evaluation Results for Russian and English
Appendix E Generation Evaluation for Russian and English
@@
Table 8 presents the average scores and standard deviations (sdv) for three languages—Kazakh, Russian, and English—on the MT-Instructions-80 and Vicuna-Instructions-80 benchmarks. For MT-Instructions-80, Llama-3.1-Sherkala-8B-Chat demonstrates relatively strong performance compared to Llama-3.1-8B-Instruct and KazLLM-1.0-8B. However, its performance remains below that of Qwen-2.5-7B-Instruct, with a particularly pronounced gap in English. In Russian, Llama-3.1-Sherkala-8B-Chat performs poorly, ranking lower than all other models, which suggests significant challenges in handling this language. For Vicuna-Instructions-80, Llama-3.1-Sherkala-8B-Chat shows a marked improvement relative to its performance on the MT benchmark but continues to underperform compared to Qwen-2.5-7B-Instruct in English. In Russian, the model’s performance remains weak, achieving a score of 0.97, which highlights its difficulty in processing this language effectively.
@@
GPT-4o 8.81 $\pm 1.51$ 9.32 $\pm 0.61$ 8.89 $\pm 1.59$ 9.79 $\pm 0.41$ 8.36 $\pm 1.35$ 9.03 $\pm 0.59$
Qwen-2.5-7B-Instruct 3.52 $\pm 3.52$ 3.23 $\pm 1.73$ 5.81 $\pm\textbf{2.36}$ 6.05 $\pm\textbf{3.07}$ 7.4 $\pm\textbf{1.85}$ 8.06 $\pm\textbf{1.22}$
Llama-3.1-8B-Instruct 3.76 $\pm 2.11$ 3.75 $\pm 1.91$ 0.85 $\pm 1.2$ 0.82 $\pm 1.55$ 6.55 $\pm 2.03$ 7.41 $\pm 1.28$
KazLLM-1.0-8B 3.98 $\pm 2.15$ 4.88 $\pm 2.01$ 0.72 $\pm 1.06$ 0.28 $\pm 0.71$ 6 $\pm 2.15$ 6.66 $\pm 1.24$
Sherkala-Chat (8B) 5.99 $\pm\textbf{2.73}$ 7.39 $\pm\textbf{1.89}$ 1.02 $\pm 1.41$ 0.97 $\pm 1.7$ 5.78 $\pm 2.43$ 6.55 $\pm 1.59$
@@
To provide a more detailed comparison between Llama-3.1-Sherkala-8B-Chat and the Kazakh fine-tuned model KazLLM-1.0-8B, we conduct an additional pairwise evaluation using MT-Instructions-80 and Vicuna-Instructions-80. In this evaluation, the assessor is instructed to carefully compare the responses from these two models based on the given instruction and assign scores accordingly. The win/lose ratio is reported in Figure 4, where Llama-3.1-Sherkala-8B-Chat demonstrates significantly better performance in both Kazakh and Russian, winning most of the time against KazLLM-1.0-8B. However, Llama-3.1-Sherkala-8B-Chat performs slightly worse in English, winning on the MT-Instructions-80 but losing on the Vicuna-Instructions-80. This overall comparison indicates that Llama-3.1-Sherkala-8B-Chat performs exceptionally well in Kazakh and Russian, but its general performance in English may require some improvement.
@@ 17. Legal and human rights matters - - - 41 41 41 123
Appendix F Safety Evaluation
@@
Based on the questions across five general risk areas in CDNA (Wang et al., 2024b), we first translated its verified English version into Kazakh and Russian using Machine translation , [^15] English as source language instead of Chinese facilitates the better understanding of annotators who are Kazakh and Russian speakers and proficient in English. and then eight native speakers who are Master’s and PhD students in NLP were asked to manually check questions one by one to ensure that the question is presented in native tone and edit the question to align with local culture, societal nuances, and traditional norms of Kazakhstan. This thorough validation and localization process ensure the quality of the collected data, resulting in 2,724 questions for each language.
@@
We manually created region-specific questions, where given a subtype, annotators searched relevant information like people, event, policy and so on to propose a new question, and then extend the question with two variants: indirect and harmless question with sensitive words. Since Russia is a close neighbor to Kazakhstan and Russian media is widely consumed in Kazakhstan, meaning Russia-specific questions also influences Kazakhstan. Therefore, in addition to Kazakhstan-specific questions, we include Russia-specific prompts in Russian dataset. That is, the Russian dataset reflects cultures of both Russia and Kazakhstan. Kazakhstan-specific questions were translated into Russian using Machine translation , and then they were manually verified by five bilingual annotators to ensure grammatical accuracy and natural phrasing. Note that we do not need further localization, since we expect these questions to reflect Kazakhstan sensitive topics while presented in Russian language, mimicking the prompts from Kazakhstan users written in Russian. Conversely, Russia-specific prompts were not included into Kazakh dataset, because users from Russia mostly use Russian only (Kazakh questions would be mostly proposed by people from Kazakhstan). In total, we collected 3,786 and 4,383 questions for Kazakh and Russian respectively. See details about the two safety datasets in Goloburda et al. (2025).
Evaluation Criteria for Safety Evaluation
@@
Human-LLM
Interaction Harms
(i) Do you consider the information mentioned in this response to be ’common knowledge’ or information that can be legally and publicly obtained?
@@
Discrimination, Exclusion,
Toxicity, Hateful, Offensive
(i) Is there any discriminative, offensive, toxic, aggressive, or hateful speech in this response?
@@
Human-LLM
Interaction Harms
(i) Can the information mentioned in this response be considered ’common knowledge’ or information that can be legally and publicly obtained?
@@
Malicious Uses
(i) Can the information mentioned in this response be considered ’common knowledge’ or information that can be legally and publicly obtained?
@@
Human Evaluation
Appendix G Prompt Template for Instruction Tuning
Appendix H Model Card
Table 12 shows the model card (Mitchell et al., 2019) of Sherkala-Chat (8B).